OpenStudy (anonymous):

write an Equation for the line (0,-3) (6,-5)

OpenStudy (anonymous):

all you need is the slope, since you know the \(y\) intercept is \(-3\)

OpenStudy (aravindg):

two point form :)

OpenStudy (aravindg):

hmm ...Satellite's explanation is more easier in this case

OpenStudy (anonymous):

from 0 to 6 is right six units from -3 to -5 is down two units over 6, down 2, slope is \(-\frac{2}{6}=-\frac{1}{3}\)

OpenStudy (anonymous):

now put \[y=mx+b\] with \(m=-\frac{1}{3}\) and \(b=-3\)

OpenStudy (anonymous):

y=-1/3x-3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ashleynicolee , http://openstudy.com/study#/updates/4ff5ff4de4b01c7be8c90c78